Calculate the amount of heat released by a given reaction, causing the temperature of 500 mL water (density = 1 g/mL) to increase by 278.35 K (5.2°C)

1
Expert's answer
2022-03-25T15:17:01-0400

q = mC∆T

q = heat

m = mass = 50 g NaCl + 100 g AgNO3 = 150 g

C = specific heat = 4.20 J/g/deg

∆T = change in temperature = 25.2 - 24.1 = 1.1 degrees 

q = (150.0 g)(4.20 J/g/deg)(1.1 deg) = 693 joules